OpenStudy (anonymous):

ΔABC is a scalene triangle. m∠A = (2x + 8)°, m∠B = (2x - 2)°, and m∠C = (3x - 8)°. Find the measure of the largest angle.

OpenStudy (anonymous):

2x+8 3x-8 2x-2 add=7x-2=180 deg in a triangle so x=182/7=26 deg A=2*26+8=60 B=2*26-2=50 C=3*26-8=70 check sum=180deg